1+tanA=√2 sin(45+A)/cosA=√2 cos(45-A)/cosA how can 1+tanA be written as above.???


show the derivations of both 1+tanA=√2 sin(45+A)/cosA and 1+tan A =√2 cos(45-A)/cosA


need it fast.and derivation is really important.

Grade:10

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Its derivation is very simple...

1+tanA=1+sinA/cosA=sinA+cosA/cosA

multiply and divide by root 2

 we get

√2 sin(45+A)/cosA=√2 cos(45-A)/cosA

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

vikas askiitian expert
13 years ago

1+ tanA = (sinA + cosA)/cosA                           (multiply divide  by 21/2)

 

1+ tanA = 21/2[sinA(1/21/2) + cosA(1/21/2) ]/cosA                  .................1

   sinpi/4 = cospi/4  =   1/21/2

now , eq  1 can be rewritten as

1+tanA = 21/2 [sinAsinpi/4+cosAcospi/4] /cosA     or  21/2[sinAcospi/4+cosAsinpi/4]/cosA

 

now we have , sinacosb+cosasinb=sina+b    &      cosacosb+sinasinb=cosa-b

using these

1+tanA = 21/2[cos(pi/4-A)]/cosA = 21/2[sin(pi/4+A)]/cosA

(pi/4 =45)

hence proved